OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The average of two test scores is 79. If one test score is eleven less than the other, what are the two test scores?

OpenStudy (cwtan):

x+y=79 x=11+y 11+2y=79 2y=68 y=34 x=79-34 =45

OpenStudy (anonymous):

thanks!! I'm about to work it out by myself to see if I come up with the same answer :)